# A dynamical approach to studying the Lee-Yang zeros for the Potts Model on the Cayley Tree

### v1

@@ Abstract.
Let $Z_{n}(z,t)$ denote the partition function of the $q$-state Potts Model on the rooted binary Cayley tree of depth $n$. Here, $z={\rm e}^{-h/T}$ and $t={\rm e}^{-J/T}$ with $h$ denoting an externally applied magnetic field, $T$ the temperature, and $J$ a coupling constant. One can interpret $z$ as a “magnetic field-like” variable and $t$ as a “temperature-like” variable. Physical values $h\in\mathbb{R}$, $T>0$, and $J\in\mathbb{R}$ correspond to $t\in(0,\infty)$ and $z\in(0,\infty)$. For any fixed $t_{0}\in(0,\infty)$ and fixed $n\in\mathbb{N}$ we consider the complex zeros of $Z_{n}(z,t_{0})$ and how they accumulate on the ray $(0,\infty)$ of physical values for $z$ as $n\rightarrow\infty$. In the ferromagnetic case ($J>0$ or equivalently $t\in(0,1)$) these Lee-Yang zeros accumulate to at most one point on $(0,\infty)$ which we describe using explicit formulae. In the antiferromagnetic case $(J<0$ or equivalently $t\in(1,\infty)$) these Lee-Yang zeros accumulate to finitely many points of $(0,\infty)$, which we again describe with explicit formulae. The same results hold for the unrooted Cayley tree of branching number two.
@@ 1.1. Lee-Yang zeros for the Ising Model
In the Ising model, one magnetic particle (an electron, for example) is at each vertex, and two particles interact if and only if they are connected by an edge. Each particle is assigned a magnetic moment, called spin, which is represented in the model by the discrete variable $\sigma(i)\in\{-1,+1\}$ which describes the spin at vertex $i$. For each spin configuration, $\sigma:V_{n}\rightarrow\{\pm 1\}$ define the Hamiltonian (energy) of $\sigma$ by
@@
For a fixed temperature $T>0$, the Gibbs-Boltzman weight of the spin configuration $\sigma$ is given by
@@
(We set the Boltzman constant $k_{B}=1$.) Summing over all the possible spin configurations gives the partition function
@@
The probability $P_{n}(\sigma)$ of the system being in the state $\sigma$ is proportional to $W_{n}(\sigma)$. Thus,
@@ 1.4. Binary Cayley Tree
The $n^{th}$-level rooted binary Cayley tree is a tree (a simple, undirected, connected, and finite graph in which any two vertices are connected by exactly one path) in which one vertex, called the root, is of degree two with all leaves (vertices of degree one) at a distance (minimum number of edges to connect) $n$ from the root, and all the other vertices are of degree three. The $n^{th}$-level unrooted binary Cayley tree is a tree in which all vertices have degree $3$ or $1$ and for which there exists a unique vertex of degree $3$ such that all of the leaves are of distance $n$ from that specified vertex. The remainder of this paper we will denote the rooted binary Cayley Tree of level $n$ by $\Gamma_{n}$ and the unrooted binary Cayley Tree of level $n$ by $\hat{\Gamma}_{n}$. We will consider the $q$-state Potts Model on these two families of graphs, as $n$ approaches infinity, for the remainder of this paper.
@@
The left side of Figure 2 shows the Lee-Yang zeros for the Ising model on the binary Cayley tree when $n=5$ and when $t=0.0625$. Note that the zeros appear to lie perfectly on the unit circle $\mid z\mid=1$ as described by the Lee-Yang circle theorem. Moreover, there exists a critical temperature $t_{\rm crit}=1/3>0$ such that when $t<t_{\rm crit}$ the Lee-Yang zeros accumulate to $1$ on the positive real axis and such that when $t>t_{\rm crit}$, the Lee-Yang zeros stay away from the positive real axis (right side of Figure 2). (For a rigorous justification of this phenomenon, see [13, Theorem A].)
The situation is dramatically different for the $3$-state Potts model on the binary Cayley tree because the Lee-Yang zeros no longer lie on the unit circle. In fact, the Lee-Yang zeros of the $5^{th}$ rooted binary Cayley tree for the $3$-state Potts model seem to lie inside the unit circle (left side of Figure 3). Furthermore, there is a critical temperature $t_{\rm crit}=\frac{1+\sqrt{73}}{36}\approx 0.265$ such that for any $t\leq t_{\rm crit}$, these zeros accumulate to a point on the positive real axis and such that when $t>t_{c}$, the Lee-Yang zeros stay away from the positive real axis (right side of Figure 3). (This will be rigorously justified in Theorem A.)
Figure 2. The Lee-Yang zeros for the $5^{th}$ rooted Cayley Tree with branching number 22 (Ising model). Left: $t=0.0625<t_{c}=1/3$. Right: $t=0.5>t_{c}=1/3$.
Figure 3. The Lee-Yang zeros for the $5^{th}$ rooted Cayley Tree with branching number 22 for 33-state Potts model. Left: $t=0.0625<t_{c}=\frac{1+\sqrt{73}}{36}\approx 0.265$. Right: $t=0.5>t_{c}=\frac{1+\sqrt{73}}{36}\approx 0.265$.
@@ 1.6. Main Results
Our results consider the sets
@@
Figure 4. Graph of $z=z_{c}(t,3)$ (Theorem A). For $0\leq t\leq 1/4$ we have $z_{c}(t,3)=1$ (blue) and for $1/4\leq t\leq t_{2}(q)\approx 0.2651$ we have $z_{c}(t,3)=\mathcal{Z}_{q}(t)$ (red).
@@
    the Lee-Yang zeros of the $q$-state Potts model on the $n^{th}$ level rooted Cayley tree with branching number two are the solutions $z$ to
@@
    the Lee-Yang zeros of the $q$-state Potts model on the $n^{th}$ level unrooted Cayley tree with branching number two are the solutions $z$ to
@@
In their 1976 paper [40], Wang and Wu considered the ferromagentic $q$-state Potts Model on the Cayley tree and they computed the temperature $T_{c}(q)>0$ such that the zero magnetic field susceptibility diverges for $T<T_{c}(q)$ but converges for $T>T_{c}(q)$. (See also [42, Section I.E].) As they explain, this proves that the limiting free energy $f(T,h)$ is non-analytic as $h$ is varied about 0, at least in the range $0<T<T_{c}(q)$. Expressed in our notation, and specialized to the binary Cayley tree, their result is
@@ Acknowledgments
We are very grateful to Arnaud Cheritat for making the computer images of the active locus shown in Figures 7 through 10 and also for describing his algorithm to us. We also thank Suzanne and Brian Boyd for their help using the Dynamics Explorer computer software to help us discover the statements proved in this paper. We are very grateful to Pavel Bleher, Bruce Kitchens, and Robert Shrock for their many helpful suggestions. We thank Thomas Gauthier for his comments about the active/passive dichotomy and we thank Krishna Kalidindi for his comments on Section 2.1. Both authors were supported by NSF grant DMS-2154414.
@@
The proofs of Theorem A-C are based on the recursive formula in Theorem D and some ideas in complex dynamics, which we present in Section 2.1, below. Section 2.2 concerns application of the methods from Section 2.1 to the Renormalization mapping $R_{z,t,q}(w)$. The ideas presented in Sections 2.1 and 2.2 allow us to plot high-quality computer-generated images of the accumulation set of Lee-Yang zeros of the Potts model on the $n^{th}$-level binary Cayley tree with branching number two as $n\rightarrow\infty$. This will be done in Section 2.3.
@@
We refer the reader to [33] for general background on holomorphic dynamics and to [10, Section V] for a similar discussion of the topics presented below that is also written for an audience from statistical physics.
@@
See [15, Chapter VII] for the definition of normal family.
@@
Let $\ell$ be a natural number. Then, marked point $a(\lambda)$ is passive for $f_{\lambda}(z)$ at $\lambda=\lambda_{0}$ if and only if it is passive for the $\ell$-th iterate $f^{\ell}_{\lambda}(z)$ at $\lambda=\lambda_{0}$.
@@
Let $z_{rep}(\lambda_{0})$ be a repelling fixed point of $f_{\lambda_{0}}$. Because $f_{\lambda}$ varies holomorphically with $\lambda$, we can find a holomorphic map $z_{rep}(\lambda)$ defined in a neighborhood $U$ of $\lambda_{0}$ such that $z_{rep}(\lambda)$ is a repelling fixed point of $f_{\lambda}$ for all $\lambda\in U$. Let $n_{0}\in\mathbb{N}$. We say $f_{\lambda_{0}}^{n_{0}}$ maps $a(\lambda_{0})$ non-persistently onto the repelling fixed point $z_{rep}(\lambda_{0})$ if
@@
The definition that $f_{\lambda_{0}}$ maps the marked point $a(\lambda_{0})$ non-persistently onto a repelling periodic point of period $p>1$ is defined analogously by replacing $f_{\lambda_{0}}$ with $f_{\lambda_{0}}^{p}$.
@@
Even though this lemma is well-known in complex dynamics, for the convenience of the reader we include a sketch of the proof. Those who wish to see a more sophisticated proof can consider [22, Lemma 3.1(2)].
@@
It follows from the definition of normal families that we can assume without loss of generality that $n_{0}=0$. By the Implicit Function Theorem and conjugation of $f_{\lambda}$ by a suitable holomorphically varying Möbius transformation we can also assume without loss of generality that that $z=0$ is a repelling fixed point of $f_{\lambda}(z)$ for all $\lambda$ in some small open neighborhood $U\subset\Lambda$ of $\lambda_{0}$. If we let $r>0$ be sufficiently small then for each $\lambda\in U$ the topological annulus
@@
Our assumption that $f_{\lambda}^{n_{0}}$ maps $a(\lambda_{0})$ non-persistently onto the repelling fixed point and our assumption that $n_{0}=0$ implies that $a(\lambda_{0})=0$ and $a(\lambda)\neq 0$ for all $\lambda\neq\lambda_{0}$ chosen sufficiently close to $\lambda_{0}$. In particular for any $\lambda$ sufficiently close to $\lambda_{0}$ there exists a natural number $\ell(\lambda)=k(\lambda,a(\lambda))$ such that $f_{\lambda}^{\ell(\lambda)}(a(\lambda))\in A_{\lambda}$. For each $\lambda\in U$ and any $n\in\mathbb{N}$ we have that $f^{-n}_{\lambda}(A_{\lambda})$ is also a topological annulus surrounding 0. Therefore, it follows from the intermediate value theorem (or other basic topological considerations) that for any $n\in\mathbb{N}$ there exists $\lambda_{n}\in U$ such that $\ell(\lambda_{n})=n$.
We now suppose for contradiction that $\lambda_{0}$ is a passive parameter. Restricting $U$ to a smaller neighborhood of $\lambda_{0}$, if necessary, we can suppose that $\lambda\mapsto f_{\lambda}^{n}(a(\lambda))$ is a normal family on $U$. In particular this implies that there is some subsequence $(n_{j})_{j}$ of the sequence of all natural numbers such that $f_{\lambda}^{n_{j}}(a(\lambda))$ converges uniformly on compact subsets of $U$ to a holomorphic function $g(\lambda)$. Note that for $\lambda=\lambda_{0}$ we have that $f_{\lambda_{0}}^{n_{j}}(a(\lambda_{0}))=f_{\lambda_{0}}^{n_{j}}(0)=0$ for all $j$ implying that $g(0)=0$. On the other hand, as explained in the previous paragraph, for each $j$ there exists $\lambda_{j}\in U$ such that $\ell(\lambda_{j})=n_{j}$. Moreover we must have $\lambda_{j}\rightarrow 0$ as $j\rightarrow\infty$ because for all $\lambda$ outside of a given neighborhood of $\lambda_{0}$ we have that $|a(\lambda_{0})|$ is uniformly bounded below, implying that $\ell(\lambda)$ is uniformly bounded above for such parameters. Because the convergence of $f_{\lambda}^{n_{j}}(a(\lambda))$ to $g(\lambda)$ is uniform on compact subsets of $U$ and because $f_{\lambda_{k}}^{n_{k}}(\lambda_{k})\in A_{\lambda}$ for each index $k$ we therefore have
$|g(0)|=\lim_{k\rightarrow\infty}|f_{\lambda_{k}}^{n_{k}}(\lambda_{k})|\geq r>0.$
@@
Now suppose that $a(\lambda_{0})$ is in the basin of attraction of an attracting fixed point for $f_{\lambda_{0}}(z)$. As in the previous case, we can assume without loss of generality that $z=0$ is an attracting fixed point for $f_{\lambda}(z)$ for all $\lambda$ in some sufficiently small neighborhood of $\lambda_{0}$. Restricting $U$ further if necessary, there is a uniform choice of $r>0$ such that for all $\lambda\in U$ the disc $\mathbb{D}_{r}(0)$ is in the basin of attraction of 0 under iteration of $f_{\lambda}(z)$. By hypothesis there is a natural number $n_{0}$ such that $|f_{\lambda_{0}}^{n_{0}}(a(\lambda))|<r/2$. By continuity we will then have that $|f_{\lambda_{0}}^{n_{0}}(a(\lambda))|<3r/4$ for all $\lambda$ is some open neighborhood $V\subset U$ of $\lambda_{0}$. It is then immediate that the sequence of functions $\lambda\mapsto f_{\lambda}^{n}(a(\lambda))$ converges uniformly to $g(\lambda)\equiv 0$ on all of $V$. Therefore $\lambda_{0}$ is a passive parameter. ∎
@@
In addition to the original source [30], we also refer the reader to [10, Lemma V.2] for a proof of Lemma 4
@@
Lemma 5 gives that $\mathcal{A}(t,q)\subset\mathcal{B}(t,q)$ and $\mathcal{A}(t,q)\subset\hat{\mathcal{B}}(t,q)$. Although we are primarily interested in the sets $\mathcal{B}(t,q)$ and $\hat{\mathcal{B}}(t,q)$, the active locus $\mathcal{A}(t,q)$ for the marked point $a(z)=z$ under the holomorphic family $R_{z,t,q}(w)$ can be studied dynamically. Moreover, we can also use the computer to make plans of $\mathcal{A}(t,q)$ that are far more detailed than the plots of Lee-Yang zeros given in Figure 3. While the plots of $\mathcal{A}(t,q)$ may be missing some points of $\mathcal{B}(t,q)$ or $\hat{\mathcal{B}}(t,q)$ they still may be informative about those two sets of Lee-Yang zeros.
@@
It is difficult to numerically compute the Lee-Yang zeros of the partition function when $t>1$. Therefore, we did not include any images analogous to Figure 3 when $t>1$ (the antiferromagnetic case). Instead, we include the plots of the active locus for the marked point $a(z)=z$ for $R_{z,t}(w)$ at $t=6$ in Figure 9 and Figure 10. The Lee-Yang zeros for $\Gamma_{n}$ accumulate to this active locus as $n\rightarrow\infty$.
@@
    If Steps 1 and 2 both fail and if $m$ has reached the threshold, $m_{0}$ then the program stops and colors the pixel red, indicating that it could not decide if $a(\lambda)$ should be interpreted as being active or passive at the parameter $\lambda_{0}$.

### v2

@@ Abstract.
Let $Z_{n}(z,t)$ denote the partition function of the $q$-state Potts Model on the rooted binary Cayley tree of depth $n$. Here, $z={\rm e}^{-h/T}$ and $t={\rm e}^{-J/T}$ with $h$ denoting an externally applied magnetic field, $T$ the temperature, and $J$ a coupling constant. One can interpret $z$ as a “magnetic field-like” variable and $t$ as a “temperature-like” variable. Physical values $h\in\mathbb{R}$, $T>0$, and $J\in\mathbb{R}$ correspond to $t\in(0,\infty)$ and $z\in(0,\infty)$. For any fixed $t_{0}\in(0,\infty)$ and fixed $n\in\mathbb{N}$ we consider the complex zeros of $Z_{n}(z,t_{0})$ and how they accumulate on the ray $(0,\infty)$ of physical values for $z$ as $n\rightarrow\infty$. In the ferromagnetic case ($J>0$ or equivalently $t\in(0,1)$) these Lee-Yang zeros accumulate to at most one point on $(0,\infty)$ which we describe using explicit formulae. In the antiferromagnetic case $(J<0$ or equivalently $t\in(1,\infty)$) these Lee-Yang zeros accumulate to at most two points of $(0,\infty)$, which we again describe with explicit formulae. The same results hold for the unrooted Cayley tree of branching number two.
@@ 1. Introduction
Throughout this paper we will use the convention that the natural numbers are all integers greater than or equal to $1$, i.e. $\mathbb{N}=\{1,2,\ldots\}$. Given $k\in\mathbb{N}$ we let $\mathbb{N}_{\geq k}:=\{n\in\mathbb{N}\,:\,n\geq k\}$.
@@ 1.1. Lee-Yang zeros for the Ising Model
In the Ising model, one magnetic particle (an electron, for example) is at each vertex, and two particles interact if and only if they are connected by an edge. Each particle is assigned a magnetic moment, called spin, which is represented in the model by the discrete variable $\sigma(i)\in\{-1,+1\}$ which describes the spin at vertex $i$. For each spin configuration, $\sigma:V_{n}\rightarrow\{\pm 1\}$, define the Hamiltonian (energy) of $\sigma$ by
@@
For a fixed temperature $T>0$, the Gibbs-Boltzmann weight of the spin configuration $\sigma$ is given by
@@
(We set the Boltzmann constant $k_{B}=1$.) Summing over all the possible spin configurations gives the partition function
@@
By the Gibbs-Boltzmann hypothesis, the probability $P_{n}(\sigma)$ of the system being in the state $\sigma$ is proportional to $W_{n}(\sigma)$. Thus,
@@ 1.4. Binary Cayley Tree
The $n^{th}$-level rooted binary Cayley tree is a tree (a simple, undirected, connected, and finite graph in which any two vertices are connected by exactly one path) in which one vertex, called the root, is of degree two with all leaves (vertices of degree one) at a distance (minimum number of edges to connect) $n$ from the root, and all the other vertices are of degree three. The $n^{th}$-level unrooted binary Cayley tree is a tree in which all vertices have degree $3$ or $1$ and for which there exists a unique vertex of degree $3$ such that all of the leaves are of distance $n$ from that specified vertex. In the remainder of this paper we will denote the rooted binary Cayley Tree of level $n$ by $\Gamma_{n}$ and the unrooted binary Cayley Tree of level $n$ by $\hat{\Gamma}_{n}$. We will consider the $q$-state Potts Model on these two families of graphs, as $n$ approaches infinity, for the rest of this paper.
@@
Remark 1.
It is important to note that many classical treatments of the Ising Model on the Cayley Tree consider the thermodynamical properties associated only with vertices “deep” in the lattice (e.g. the root vertex); see [5, Ch. 4] and the references therein. The term “Bethe Lattice” is typically used to describe such considerations. Instead, we treat all vertices equally, studying the “bulk” behavior of the lattice, and thus we follow the standard convention of referring to our work as being on the Cayley Tree.
@@
The left side of Figure 2 shows the Lee-Yang zeros for the Ising model on the binary Cayley tree when $n=5$ and when $t=0.0625$. Note that the zeros appear to lie perfectly on the unit circle $\mid z\mid=1$ as described by the Lee-Yang circle theorem. Moreover, there exists a critical temperature $t_{\rm crit}=1/3>0$ such that when $t\leq t_{\rm crit}$ the Lee-Yang zeros accumulate to $1$ on the positive real axis and such that when $t>t_{\rm crit}$, the Lee-Yang zeros stay away from the positive real axis (right side of Figure 2). (For a rigorous justification of this phenomenon, see [14, Theorem A].)
The situation is dramatically different for the $3$-state Potts model on the binary Cayley tree because the Lee-Yang zeros no longer lie on the unit circle. In fact, the Lee-Yang zeros of the $5^{th}$ rooted binary Cayley tree for the $3$-state Potts model seem to lie inside the unit circle (left side of Figure 3). However, like for the Ising Model, there is a critical temperature $t_{\rm crit}=\frac{1+\sqrt{73}}{36}\approx 0.265$ such that for any $t\leq t_{\rm crit}$, these zeros accumulate to a point on the positive real axis and such that when $t>t_{\rm crit}$, the Lee-Yang zeros stay away from the positive real axis (right side of Figure 3). This will be rigorously justified in Theorem A. (The reader may also compare this description with a conjectural description of the accumulation locus of the Lee-Yang zeros described in Remark 7.)
Figure 2. The Lee-Yang zeros for the $5^{th}$ rooted binary Cayley Tree for the Ising model. Left: $t=0.0625<t_{\rm crit}=1/3$. Right: $t=0.5>t_{\rm crit}=1/3$.
Figure 3. The Lee-Yang zeros for the $5^{th}$ rooted binary Cayley Tree for the $3$-state Potts model. Left: $t=0.0625<t_{\rm crit}=\frac{1+\sqrt{73}}{36}\approx 0.265$. Right: $t=0.5>t_{\rm crit}=\frac{1+\sqrt{73}}{36}\approx 0.265$.
@@ 1.6. Main Results
For fixed choices of $t\in[0,\infty]$ and $q\in\mathbb{N}_{\geq 2}$ our results consider the sets
@@
Figure 4. Graph of $z=z_{c}(t,3)$ (Theorem A). For $0\leq t\leq 1/4$ we have $z_{c}(t,3)=1$ (blue) and for $1/4<t\leq t_{2}(q)\approx 0.2651$ we have $z_{c}(t,3)=\mathcal{Z}_{q}(t)<1$ (red).
@@
    the Lee-Yang zeros of the $q$-state Potts model on the $n^{th}$ level rooted binary Cayley tree are the solutions $z$ to
@@
    the Lee-Yang zeros of the $q$-state Potts model on the $n^{th}$ level unrooted binary Cayley tree are the solutions $z$ to
@@
In their 1976 paper [41], Wang and Wu considered the ferromagnetic $q$-state Potts Model on the Cayley tree and they computed the temperature $T_{c}(q)>0$ such that the zero magnetic field susceptibility diverges for $T<T_{c}(q)$ but converges for $T>T_{c}(q)$. (See also [43, Section I.E].) As they explain, this proves that the limiting free energy $f(T,h)$ is non-analytic as $h$ is varied about 0, at least in the range $0<T<T_{c}(q)$. Expressed in our notation, and specialized to the binary Cayley tree, their result is
@@ Acknowledgments
We are very grateful to Arnaud Cheritat for making the computer images of the active locus shown in Figures 7 through 10 and also for describing his algorithm to us. We also thank Suzanne and Brian Boyd for their help using the Dynamics Explorer computer software to help us discover the statements proved in this paper. We are very grateful to Pavel Bleher, Bruce Kitchens, and Robert Shrock for their many helpful suggestions. We thank Thomas Gauthier for his comments about the active/passive dichotomy and we thank Krishna Chaitanya Kalidindi, Andres Quintero, and Yapeng Zhao for their comments on the writing. Both authors were supported by NSF grant DMS-2154414.
@@
The proofs of Theorem A-C are based on the recursive formula in Theorem D and some ideas in complex dynamics, which we present in Section 2.1, below. Section 2.2 concerns application of the methods from Section 2.1 to the Renormalization mapping $R_{z,t,q}(w)$. The ideas presented in Sections 2.1 and 2.2 allow us to plot high-quality computer-generated images of the accumulation set of Lee-Yang zeros of the Potts model on the $n^{th}$-level binary Cayley tree as $n\rightarrow\infty$. This will be done in Section 2.3.
@@
We refer the reader to [34] for general background on holomorphic dynamics and to [11, Section V] for a similar discussion of the topics presented below that is also written for an audience from statistical physics.
@@
See [16, Chapter VII] for the definition of normal family.
@@
Let $\ell\geq 2$ be a natural number. Then, marked point $a(\lambda)$ is passive for $f_{\lambda}(z)$ at $\lambda=\lambda_{0}$ if and only if it is passive for the $\ell$-th iterate $f^{\ell}_{\lambda}(z)$ at $\lambda=\lambda_{0}$.
@@
Let $z_{rep}(\lambda_{0})$ be a repelling fixed point of $f_{\lambda_{0}}$. Because $f_{\lambda}$ varies holomorphically with $\lambda$, we can find a holomorphic map $z_{rep}(\lambda)$ defined in a neighborhood $U$ of $\lambda_{0}$ such that $z_{rep}(\lambda)$ is a repelling fixed point of $f_{\lambda}$ for all $\lambda\in U$. Let $n_{0}\in\mathbb{N}\cup\{0\}$. We say $f_{\lambda_{0}}^{n_{0}}$ maps $a(\lambda_{0})$ non-persistently onto the repelling fixed point $z_{rep}(\lambda_{0})$ if
@@
(Here, $f_{\lambda}^{0}$ is interpretted as the identity map.) The definition that $f_{\lambda_{0}}$ maps the marked point $a(\lambda_{0})$ non-persistently onto a repelling periodic point of period $p>1$ is defined analogously by replacing $f_{\lambda_{0}}$ with $f_{\lambda_{0}}^{p}$.
@@
Even though this lemma is well-known in complex dynamics, for the convenience of the reader we include a sketch of the proof. Those who wish to see an alternative proof can consider [23, Lemma 3.1(2)].
@@
It follows from the definition of normal families that we can assume without loss of generality that $n_{0}=0$. By the Implicit Function Theorem and conjugation of $f_{\lambda}$ by a suitable holomorphically varying Möbius transformation we can also assume without loss of generality that $z=0$ is a repelling fixed point of $f_{\lambda}(z)$ for all $\lambda$ in some small open neighborhood $U\subset\Lambda$ of $\lambda_{0}$. If we let $r>0$ be sufficiently small then for each $\lambda\in U$ the topological annulus
@@
Our assumption that $f_{\lambda}^{n_{0}}$ maps $a(\lambda_{0})$ non-persistently onto the repelling fixed point and our assumption that $n_{0}=0$ implies that $a(\lambda_{0})=0$ and $a(\lambda)\neq 0$ for all $\lambda\neq\lambda_{0}$ chosen sufficiently close to $\lambda_{0}$. In particular for any $\lambda$ sufficiently close to $\lambda_{0}$ there exists a natural number $\ell(\lambda)=k(\lambda,a(\lambda))$ such that $f_{\lambda}^{\ell(\lambda)}(a(\lambda))\in A_{\lambda}$. For each $\lambda\in U$ and any $n\in\mathbb{N}$ we have that $f^{-n}_{\lambda}(A_{\lambda})$ is also a topological annulus surrounding 0. Therefore, it follows from the intermediate value theorem applied to $|a(\lambda)|$ (or other basic topological considerations) that for any $n\in\mathbb{N}$ there exists $\lambda_{n}\in U$ such that $\ell(\lambda_{n})=n$.
We now suppose for contradiction that $\lambda_{0}$ is a passive parameter. Restricting $U$ to a smaller neighborhood of $\lambda_{0}$, if necessary, we can suppose that $\lambda\mapsto f_{\lambda}^{n}(a(\lambda))$ is a normal family on $U$. In particular this implies that there is some subsequence $(n_{j})_{j}$ of the sequence of all natural numbers such that $f_{\lambda}^{n_{j}}(a(\lambda))$ converges uniformly on compact subsets of $U$ to a holomorphic function $g(\lambda)$. Note that for $\lambda=\lambda_{0}$ we have that $f_{\lambda_{0}}^{n_{j}}(a(\lambda_{0}))=f_{\lambda_{0}}^{n_{j}}(0)=0$ for all $j$ implying that $g(0)=0$. On the other hand, as explained in the previous paragraph, for each $j$ there exists $\lambda_{j}\in U$ such that $\ell(\lambda_{j})=n_{j}$. Moreover we must have $\lambda_{j}\rightarrow 0$ as $j\rightarrow\infty$ because for all $\lambda$ outside of a given neighborhood of $\lambda_{0}$ we have that $|a(\lambda)|$ is uniformly bounded below, implying that $\ell(\lambda)$ is uniformly bounded above for such parameters. Because the convergence of $f_{\lambda}^{n_{j}}(a(\lambda))$ to $g(\lambda)$ is uniform on compact subsets of $U$ and because $f_{\lambda_{k}}^{n_{k}}(a(\lambda_{k}))\in A_{\lambda}$ for each index $k$ we therefore have
$|g(0)|=\lim_{k\rightarrow\infty}|f_{\lambda_{k}}^{n_{k}}(a(\lambda_{k}))|\geq r>0.$
@@
Now suppose that $a(\lambda_{0})$ is in the basin of attraction of an attracting fixed point for $f_{\lambda_{0}}(z)$. As in the previous case, we can assume without loss of generality that $z=0$ is an attracting fixed point for $f_{\lambda}(z)$ for all $\lambda$ in some sufficiently small neighborhood $U$ of $\lambda_{0}$. Restricting $U$ further if necessary, there is a uniform choice of $r>0$ such that for all $\lambda\in U$ the disc $\mathbb{D}_{r}(0)$ is in the basin of attraction of 0 under iteration of $f_{\lambda}(z)$. By hypothesis there is a natural number $n_{0}$ such that $|f_{\lambda_{0}}^{n_{0}}(a(\lambda_{0}))|<r/2$. By continuity we will then have that $|f_{\lambda_{0}}^{n_{0}}(a(\lambda))|<3r/4$ for all $\lambda$ in some open neighborhood $V\subset U$ of $\lambda_{0}$. It is then immediate that the sequence of functions $\lambda\mapsto f_{\lambda}^{n}(a(\lambda))$ converges uniformly to $g(\lambda)\equiv 0$ on all of $V$. Therefore $\lambda_{0}$ is a passive parameter. ∎
@@
In addition to the original source [31], we also refer the reader to [11, Lemma V.2] for a proof of Lemma 4
@@
Lemma 5 gives that $\mathcal{A}(t,q)\subset\mathcal{B}(t,q)$ and $\mathcal{A}(t,q)\subset\hat{\mathcal{B}}(t,q)$. Although we are primarily interested in the sets $\mathcal{B}(t,q)$ and $\hat{\mathcal{B}}(t,q)$, the active locus $\mathcal{A}(t,q)$ for the marked point $a(z)=z$ under the holomorphic family $R_{z,t,q}(w)$ can be studied dynamically. Moreover, we can also use the computer to make plots of $\mathcal{A}(t,q)$ that are far more detailed than the plots of Lee-Yang zeros given in Figure 3. While the plots of $\mathcal{A}(t,q)$ may be missing some points of $\mathcal{B}(t,q)$ or $\hat{\mathcal{B}}(t,q)$ they still may be informative about those two sets of Lee-Yang zeros.
@@
Remark 7.
We have made computer images of the active locus for $q=3$ at many values of $t\in(0,1)$. They all seem to have the following property:
  • (i)
    If $0<t\leq t_{2}(3)$ we have that $(\mathbb{C}\setminus\{0\})\setminus\mathcal{A}(t,3)$ consists of two connected components, one of which contains $z=0$ in its closure and the other which contains $z=\infty$ in its closure. See Figure 7 for an example.
  • (ii)
    If $t>t_{2}(3)$ then $(\mathbb{C}\setminus\{0\})\setminus\mathcal{A}(t,3)$ consists of a single connected component that contains both $z=0$ and $z=\infty$ in its closure. See Figure 8 for an example.
Since the active locus $\mathcal{A}(t,q)\subset\mathcal{B}(t,q)$ is expected to provide a “good approximation” to $\mathcal{B}(t,q)$, one may also expect that Properties (i) and (ii) should also hold for $\mathcal{B}(t,q)$. For the $q\geq 3$ state Potts model on the classical $\mathbb{Z}^{d}$ lattices, this behaviour for the accumulation locus of Lee-Yang zeros is expected to be true. It is a generalization of the conjectural discription for the accumulation locus of Lee-Yang zeros for the Ising Model, which is believed to be the whole unit circle for $0<t<t_{c}$ and a single arc on the circle containing z=1z=-1 for $t>t_{c}$. See, e.g. [14, Section 1.3].
For the Ising Model on the the Cayley Tree, the conjectural description for the accumulation locus of Lee-Yang zeros was proved in [14, Theorem A]. We intend to further study whether (i) and (ii) hold for the accumulation locus of Lee-Yang zeros associated to the $q\geq 3$ state Potts model on the Cayley Tree in a future work.
It is difficult to numerically compute the Lee-Yang zeros of the partition function when $t>1$. Therefore, we did not include any images analogous to Figure 3 when $t>1$ (the antiferromagnetic case). Instead, we include the plots of the active locus for the marked point $a(z)=z$ for $R_{z,t,3}(w)$ at $t=6$ in Figure 9 and Figure 10. The Lee-Yang zeros for $\Gamma_{n}$ accumulate to this active locus as $n\rightarrow\infty$.
@@
    If Steps 1 and 2 both fail and if $m$ has reached the threshold $m_{0}$ then the program stops and colors the pixel red, indicating that it could not decide if $a(\lambda)$ should be interpreted as being active or passive at the parameter $\lambda_{0}$.